# Corpus Poisoning via Approximate Greedy Gradient Descent

### v1

@@ Abstract
Dense retrievers are widely used in information retrieval and have also been successfully extended to other knowledge intensive areas such as language models, e.g., Retrieval-Augmented Generation (RAG) systems. Unfortunately, they have recently been shown to be vulnerable to corpus poisoning attacks in which a malicious user injects a small fraction of adversarial passages into the retrieval corpus to trick the system into returning these passages among the top-ranked results for a broad set of user queries. Further study is needed to understand the extent to which these attacks could limit the deployment of dense retrievers in real-world applications. In this work, we propose Approximate Greedy Gradient Descent (AGGD), a new attack on dense retrieval systems based on the widely used HotFlip method for efficiently generating adversarial passages. We demonstrate that AGGD can select a higher quality set of token-level perturbations than HotFlip by replacing its random token sampling with a more structured search. Experimentally, we show that our method achieves a high attack success rate on several datasets and using several retrievers, and can generalize to unseen queries and new domains. Notably, our method is extremely effective in attacking the ANCE retrieval model, achieving attack success rates that are 17.6% and 13.37% higher on the NQ and MS MARCO datasets, respectively, compared to HotFlip. Additionally, we demonstrate AGGD’s potential to replace HotFlip in other adversarial attacks, such as knowledge poisoning of RAG systems. [^1] Code can be find in https://github.com/JinyanSu1/AGGD
@@ 1 Introduction
  • We conduct extensive experiments to show the vulnerability of dense retrievers under AGGD. For example, when attacking the ANCE retriever, injecting just one adversarial passage (0.00003% of the NQ dataset corpus and 0.00001% of the MS MARCO dataset corpus) can achieve an attack success rate of 44.35% and 26.16% on these datasets, respectively, improving by 17.6% and 13.37% over HotFlip. The generated adversarial passage also possesses the capability to transfer to unseen queries in other domains, with AGGD achieving an attack success rate of 82.28% on the NFCorpus dataset.
@@ 2 Related Work
Adversarial Attacks in Retrieval Systems Black-hat search engine optimization, which aims to increase the exposure of certain documents through malicious manipulation, poses a threat by reducing the quality of search results and inundating users with irrelevant pages (Castillo et al., 2011; Liu et al., 2023). Previous work has shown that retrieval systems are susceptible to small perturbations: making small edits to a target passage can significantly alter its retrieval rank (Song et al., 2020; Raval & Verma, 2020; Song et al., 2022) for individual or a small set of queries. More recently, a stronger setting known as corpus poisoning attack has been studied in Zhong et al. (2023), where the attack success rate of an adversarial passage is evaluated on unseen queries rather than on targeted given queries. These attacks differ from data poisoning attacks (Long et al., 2024; Liu et al., 2023; Chen et al., 2017; Schuster et al., 2020), as adversarial passages are injected into the retrieval corpus rather than the training data of the retrievers. The retrieval model remains unchanged in a corpus poisoning attack.
Discrete Optimization Many adversarial attacks in NLP involve discrete optimization, whether in classification tasks (Wallace et al., 2019; Ebrahimi et al., 2018; Song et al., 2021), retrieval systems (Jia & Liang, 2017; Song et al., 2020; Raval & Verma, 2020; Song et al., 2022), or adversarial prompt generation (Zou et al., 2023; Shin et al., 2020; Wen et al., 2024). The objective is to find small perturbations to the input to lead the model to make erroneous predictions. Due to the discrete nature of texts, directly applying adversarial attack methods from prior computer vision research (Xiao & Wang, 2021; Tolias et al., 2019) is infeasible. Instead, many methods build upon HotFlip (Ebrahimi et al., 2018) and approximate the effect of replacing a token using gradients.
@@ 5.1 Experimental Details
Evaluation Metrics After generating the adversarial passages on the training set and injecting them into the corpus, we evaluate the effectiveness of our attack using the top-$k$ attack success rate (ASR) on test queries. Top-$k$ ASR is defined as the percentage of queries for which at least one adversarial passage is retrieved in the top-$k_{r}$ results, i.e., $\text{ASR}=\frac{1}{n_{q}}\sum_{i=1}^{n_{q}}\mathbbm{1}\{a\in\mathcal{R}_{r}(q^{\text{test}}_{i},k_{r})\}$, where $n_{q}$ is the total number of test queries and $\mathcal{R}_{r}(q_{i}^{\text{test}},k_{r})$ is the retriever that returns the top-$k_{r}$ most relevant passages for the test test query $q_{i}^{\text{test}}$. $\mathbbm{1}\{a\in r(q_{i}^{\text{test}},k_{r})\}$ is the indicator function, which equals to 1 if $a\in\mathcal{R}(q_{i}^{\text{test}},k_{r})$ and 0 otherwise. A higher ASR indicates that the model is more vulnerable to attacks, and thus, the attack is more effective. We use $k_{r}=1000$ unless specified otherwise.
@@ 5.2 Main Results
NQ AGGD 89.57(3.58) 54.53(11.29) 44.35(3.54) 1.47(0.41) 0.14(0.06)
HotFlip 88.46(0.84) 49.64(6.09) 26.75(1.45) 1.01(0.15) 0.19(0.12)
Random 74.64(1.78) 18.34(1.48) 7.75(1.88) 0.5(0.21) 0.19(0.1)
MS MARCO AGGD 75.58(5.81) 13.95(15.16) 26.16(13.3) 2.91(1.01) 2.91(1.93)
HotFlip 75.0(4.76) 13.95(9.16) 12.79(5.07) 1.74(1.01) 0.58(1.01)
Random 55.81(6.98) 1.74(3.02) 5.81(3.49) 1.16(2.01) 0.58(1.01)
Table 1: In-domain attack success rate (ASR) of AGGD on NQ and MS MARCO datasets with 5 retrievers by injecting 1 adversarial passage. (Equivalent to increasing the corpus size by 0.00003% for NQ and 0.00001% for MS MARCO.) We highlight the best performing attacking method in bold (Higher ASR indicates better attack performance). Results are from 4 random runs with standard deviation in parenthesis.
In-Domain Attack We show the in-domain attack performance of AGGD and the other two baselines (HotFlip and random perturbation) in Table 1, evaluating the injection of only one adversarial passage. Our findings are as follows: (1) The pretrained Contriever model is more vulnerable to attacks. All three attack baselines achieve the highest ASR compared to other retrieval models. Poisoning with AGGD achieves 89.57% ASR on NQ dataset and 75.58% on MS MARCO dataset, respectively. Even using Random perturbation can achieve a relatively high ASR of 74.64% on NQ dataset and 55.81% on MS MARCO dataset. (2) Besides achieving comparable results in Contriever and DPR, AGGD is extremely effective in attacking ANCE, improving over the second-best baseline by 17.6% and 13.37% on NQ and MS MARCO, respectively. The effectiveness of AGGD in attack ANCE can also be clearly observed during the training. Due space limitations, we provide training accuracy in Appendix A.2(Figure 10). (3)Supervised retriever models such as DPR are more challenging to attack. Attacking DPR-nq on NQ with only 1 adversarial passage does no better than random perturbation. (4) MS MARCO seems to be harder to attack than NQ, likely due to its larger testing corpus, and thus, lower poison rate.
@@
FiQA-2018 NQ AGGD 62.04(17.22) 3.78(1.39) 4.09(1.18) - -
HotFlip 65.28(9.06) 2.08(0.55) 2.89(0.62) 0.23(0.23) 0.15(0.11)
Random 37.38(2.76) 0.27(0.13) 0.96(0.33) 0.15(0.11) 0.31(0.11)
MS MARCO AGGD 84.57(1.97) 14.16(21.18) 11.46(1.61) 1.89(0.56) 1.97(1.24)
HotFlip 75.5(7.87) 3.24(2.06) 7.02(0.56) 1.2(0.6) 1.81(1.18)
Random 58.76(6.12) 0.23(0.32) 3.51(1.06) 0.46(0.24) 0.42(0.13)
NFCorpus NQ AGGD 51.32(7.79) 20.98(5.59) 49.54(6.21) 12.62(2.09) 16.25(0.56)
HotFlip 45.9(11.47) 13.24(2.6) 46.13(2.22) 13.85(3.57) 17.41(1.63)
Random 28.79(6.24) 8.67(2.19) 27.01(5.86) 10.29(2.03) 16.18(0.46)
MS MARCO AGGD 56.58(14.85) 22.91(18.3) 82.28(3.41) 38.16(5.66) 40.63(4.49)
HotFlip 54.49(6.28) 17.96(10.63) 65.56(4.99) 29.18(3.16) 34.98(8.55)
Random 32.97(5.71) 8.59(8.34) 55.73(7.11) 23.61(4.15) 23.14(2.81)
Quora NQ AGGD 57.36(14.93) 11.24(4.69) 24.48(3.48) 0.36(0.06) 0.27(0.14)
HotFlip 61.82(4.97) 8.09(4.42) 20.43(0.73) 0.3(0.14) 0.2(0.06)
Random 30.09(4.01) 1.03(0.1) 10.88(0.94) 0.17(0.02) 0.27(0.05)
MS MARCO AGGD 73.42(4.37) 8.25(12.74) 37.07(2.52) 3.32(1.32) 3.37(1.07)
HotFlip 69.62(8.69) 2.75(1.59) 27.92(2.31) 2.03(0.95) 2.31(0.34)
Random 42.14(7.59) 0.37(0.45) 18.39(0.85) 0.8(0.2) 0.7(0.54)
SCIDOCS NQ AGGD 21.18(7.91) 11.75(7.84) 9.55(1.72) 0.87(0.26) 0.4(0.1)
HotFlip 27.15(13.45) 14.1(4.41) 7.18(1.0) 1.18(0.4) 0.35(0.11)
Random 12.82(3.62) 0.73(0.58) 2.9(0.8) 0.57(0.37) 0.27(0.08)
MS MARCO AGGD 23.25(6.88) 14.22(20.72) 31.32(7.33) 3.92(1.7) 1.13(0.45)
HotFlip 24.98(6.23) 11.98(7.4) 21.25(4.37) 3.05(1.34) 0.85(0.32)
Random 15.02(1.6) 1.23(1.74) 10.58(5.27) 2.35(1.27) 0.23(0.08)
SciFact NQ AGGD 25.83(4.19) 1.58(0.86) 10.83(2.88) 0.17(0.17) -
HotFlip 28.83(14.18) 1.17(0.83) 6.75(0.92) 0.42(0.43) -
Random 9.58(4.44) - 1.42(0.86) - -
MS MARCO AGGD 31.83(10.62) 3.75(5.01) 48.0(7.04) 3.5(0.9) 1.58(0.49)
HotFlip 25.0(7.06) 2.5(1.89) 28.17(8.07) 2.83(0.5) 1.42(1.44)
Random 12.08(6.19) 0.17(0.17) 13.42(7.98) 1.75(1.09) 0.67(0.71)
@@
Table 2: Out-of-domain top-50 attack success rate with only 1 adversarial passage. Results are averaged over 4 random runs with standard deviation shown in parenthesis. The combinations of attack and source dataset that achieve the highest ASR on each target domain and retriever are highlighted in bold.
@@
Out-of-Domain Attack We found that the generated adversarial passages can transfer across different domains. In Table 2, we use adversarial passages generated from training set of NQ and MS MARCO and insert them into the corpora of retrieval tasks in other domains. Since the out-of-domain datasets have smaller corpora compared to NQ and MS MARCO, we use top $k_{r}=50$ ASR (rather than $k_{r}=1000$ used for in-domain attacks) to measure the effectiveness of the attack method. We found that (1) Compared to adversarial passages generated from NQ dataset, those trained from MS MARCO generally perform better in out-of-domain attacks, possibly because MS MARCO contains more training data. (2) Contriever models are still the most vulnerable to corpus poisoning attacks. For example, inserting a single adversarial passage generated by AGGD into FiQA-2018 achieves a top-$k_{r}=50$ ASR of 84.57%, and inserting it into Quora can trick the model into returning the adversarial passage in the top-50 retrieved passage for 73.42% of the queries. (3) NFCorpus is an easier domain to attack. AGGD achieves over 38% ASR even in DPR-mul and DPR-nq, which is surprising since, as shown in the in-domain attack results in Table 1, attacking DPR-mul and DPR-nq is extremely hard , even with adversarial passage trained from in-domain data.
@@ 5.3 Analysis and Ablation Study
The effect of the token length $m$
5.4 Extending AGGD to Knowledge Poisoning Attacks
@@ 6 Conclusion
7 Limitations and Broader Impact
Though we use corpus poisoning attack as the main example to deliver our attack algorithm AGGD, our proposed attack can be used for various attack scenarios, thus, we didn’t specifically investigate defenses of AGGD for corpus poison attack. Since our method assembles to HotFlip, existing defenses to these attacks should be able to defend our attack to some extend. However, as our attack is stronger than HotFlip, it remains to understand how well these defense works.

### v2

@@ Abstract
Dense retrievers are widely used in information retrieval and have also been successfully extended to other knowledge intensive areas such as language models, e.g., Retrieval-Augmented Generation (RAG) systems. Unfortunately, they have recently been shown to be vulnerable to corpus poisoning attacks in which a malicious user injects a small fraction of adversarial passages into the retrieval corpus to trick the system into returning these passages among the top-ranked results for a broad set of user queries. Further study is needed to understand the extent to which these attacks could limit the deployment of dense retrievers in real-world applications. In this work, we propose Approximate Greedy Gradient Descent (AGGD), a new attack on dense retrieval systems based on the widely used HotFlip method for efficiently generating adversarial passages. We demonstrate that AGGD can select a higher quality set of token-level perturbations than HotFlip by replacing its random token sampling with a more structured search. Experimentally, we show that our method achieves a high attack success rate on several datasets and using several retrievers, and can generalize to unseen queries and new domains. Notably, our method is extremely effective in attacking the ANCE retrieval model, achieving attack success rates that are 15.24% and 17.44% higher on the NQ and MS MARCO datasets, respectively, compared to HotFlip. Additionally, we demonstrate AGGD’s potential to replace HotFlip in other adversarial attacks, such as knowledge poisoning of RAG systems.
Corpus Poisoning via Approximate Greedy Gradient Descent
Jinyan Su1, Preslav Nakov2, Claire Cardie1 1Department of Computer Science, Cornell University 2Mohamed bin Zayed University of Artificial Intelligence {js3673,ctc9}@cornell.edu, preslav.nakov@mbzuai.ac.ae
@@ 1 Introduction
  • We conduct extensive experiments to show the vulnerability of dense retrievers under AGGD. For example, when attacking the ANCE retriever, injecting just one adversarial passage can achieve an attack success rate of 80.92% and 65.68% on these datasets, respectively, improving by 15.24% and 17.44% over HotFlip. The generated adversarial passage also possesses the capability to transfer to unseen queries in other domains.
@@ 2 Related Work
Adversarial Attacks in Retrieval Systems Black-hat search engine optimization, which aims to increase the exposure of certain documents through malicious manipulation, poses a threat by reducing the quality of search results and inundating users with irrelevant pages (Castillo et al., 2011; Liu et al., 2023). Previous work has shown that retrieval systems are susceptible to small perturbations: making small edits to a target passage can significantly alter its retrieval rank (Song et al., 2020; Raval and Verma, 2020; Song et al., 2022) for individual or a small set of queries. More recently, a stronger setting known as corpus poisoning attack has been studied in Zhong et al. (2023), where the attack success rate of an adversarial passage is evaluated on unseen queries rather than on targeted given queries. These attacks differ from data poisoning attacks (Long et al., 2024; Liu et al., 2023; Chen et al., 2017; Schuster et al., 2020), as adversarial passages are injected into the retrieval corpus rather than the training data of the retrievers. The retrieval model remains unchanged in a corpus poisoning attack.
Discrete Optimization Many adversarial attacks in NLP involve discrete optimization, whether in classification tasks (Wallace et al., 2019; Ebrahimi et al., 2018; Song et al., 2021), retrieval systems (Jia and Liang, 2017; Song et al., 2020; Raval and Verma, 2020; Song et al., 2022), or adversarial prompt generation (Zou et al., 2023; Shin et al., 2020; Wen et al., 2024). The objective is to find small perturbations to the input to lead the model to make erroneous predictions. Due to the discrete nature of texts, directly applying adversarial attack methods from prior computer vision research (Xiao and Wang, 2021; Tolias et al., 2019) is infeasible. Instead, many methods build upon HotFlip (Ebrahimi et al., 2018) and approximate the effect of replacing a token using gradients.
@@ 5.1 Experimental Details
Evaluation Metrics After generating the adversarial passages on the training set and injecting them into the corpus, we evaluate the effectiveness of our attack using the top-$k$ attack success rate (ASR) on test queries. Top-$k$ ASR is defined as the percentage of queries for which at least one adversarial passage is retrieved in the top-$k_{r}$ results, i.e., $\text{ASR}=\frac{1}{n_{q}}\sum_{i=1}^{n_{q}}\mathbbm{1}\{a\in\mathcal{R}_{r}(q^{\text{test}}_{i},k_{r},\mathcal{C}_{\text{test}})\}$, where $n_{q}$ is the total number of test queries and $\mathcal{R}_{r}(q_{i}^{\text{test}},k_{r},\mathcal{C}_{\text{test}})$ is the retriever that returns the top-$k_{r}$ most relevant passages for the test query $q_{i}^{\text{test}}$. $\mathbbm{1}\{a\in\mathcal{R}_{r}(q_{i}^{\text{test}},k_{r},\mathcal{C}_{\text{test}})\}$ is the indicator function, which equals to 1 if $a\in\mathcal{R}_{r}(q_{i}^{\text{test}},k_{r},\mathcal{C}_{\text{test}})$ and 0 otherwise. A higher ASR indicates that the model is more vulnerable to attacks, and thus, the attack is more effective. We use $k_{r}=20$ to present our result. Since ASR depends on the size the the test corpus, to make fair comparisons across different dataset, we randomly sample $|\mathcal{C}_{\text{test}}|=10,000$ passages from the overall corpus pool.
@@ 5.2 Main Results
NQ AGGD 92.5(2.68) 63.45(8.68) 80.92(4.82) 6.88(1.72) 2.19(0.98)
Hot Flip 91.08(0.9) 58.43(4.53) 65.68(2.5) 5.4(0.36) 2.03 (0.25)
Random 80.24(0.92) 32.5(2.61) 31.0(4.3) 3.7(0.93) 1.66 (0.14)
MS MARCO AGGD 85.47(3.81) 24.42(17.44) 93.6(1.01) 12.79(3.86) 5.23 (1.01)
Hot Flip 83.72(5.93) 22.67(12.01) 76.16(9.21) 9.88(1.93) 5.81(2.01)
Random 66.86(7.24) 13.95(13.46) 49.42(11.2) 6.4(2.53) 2.91 (1.93)
Table 1: In-domain attack success rate (ASR) of AGGD on NQ and MS MARCO datasets with 5 retrievers by injecting 1 adversarial passage. We highlight the best performing attacking method in bold (Higher ASR indicates better attack performance). Results are from 4 random runs with standard deviation in parenthesis.
In-Domain Attack We show the in-domain attack performance of AGGD and the other two baselines (HotFlip and random perturbation) in Table 1, evaluating the injection of only one adversarial passage. Our findings are as follows: (1) The pretrained Contriever model is more vulnerable to attacks. All three attack baselines achieve the highest ASR compared to other retrieval models. Poisoning with AGGD achieves 92.5% ASR on NQ dataset and 85.47% on MS MARCO dataset, respectively. Even using Random perturbation can achieve a relatively high ASR of 80.24% on NQ dataset and 66.86% on MS MARCO dataset. (2) Besides achieving comparable results in Contriever and DPR, AGGD is extremely effective in attacking ANCE, improving over the second-best baseline by 15.24% and 17.44% on NQ and MS MARCO, respectively. The effectiveness of AGGD in attack ANCE can also be clearly observed during the training. Due space limitations, we provide training accuracy in Appendix A.2 (Figure 10). (3) Supervised retriever models such as DPR are more challenging to attack. Attacking DPR-nq on NQ with only 1 adversarial passage are just slightly better than random perturbation.
@@
FiQA-2018 NQ AGGD 64.93(15.9) 4.94(1.68) 6.33(1.91) 0.15(0.11) 0.12(0.13)
Hot Flip 69.68(8.66) 3.05(0.6) 4.13(0.77) 0.46(0.55) 0.62(0.15)
Random 41.05(2.73) 0.35(0.13) 1.66(0.49) 0.27(0.17) 0.46(0.19)
MS MARCO AGGD 86.54(1.88) 15.66(22.54) 17.98(1.44) 3.59(1.34) 3.36(1.79)
Hot Flip 78.51(6.96) 4.59(2.89) 11.5(0.79) 2.16(1.13) 3.2(1.72)
Random 62.81(6.07) 0.42(0.57) 5.48(1.36) 0.73(0.2) 0.93(0.24)
NFCorpus NQ AGGD 46.83(8.1) 17.1(4.3) 37.69(4.2) 8.51(1.29) 12.23(0.46)
Hot Flip 42.34(10.96) 10.91(2.77) 34.68(1.71) 10.22(2.62) 12.85(0.99)
Random 24.38(5.02) 6.58(1.43) 18.34(4.09) 7.43(1.3) 11.3(0.56)
MS MARCO AGGD 52.55(13.94) 18.73(16.3) 69.81(3.57) 26.16(4.51) 28.17(4.02)
Hot Flip 49.92(5.26) 14.24(8.46) 52.55(6.86) 19.27(2.31) 23.06(7.02)
Random 30.03(6.42) 6.42(6.45) 42.26(6.23) 14.47(3.59) 14.78(1.06)
Quora NQ AGGD 73.92(12.16) 25.1(7.03) 77.72(5.57) 4.85(1.13) 6.39(2.1)
Hot Flip 78.99(3.56) 18.92(5.84) 72.37(2.13) 3.36(1.02) 7.2(1.45)
Random 49.44(4.18) 6.12(0.63) 53.09(4.89) 2.52(0.47) 6.66(1.51)
MS MARCO AGGD 86.92(2.49) 18.83(21.47) 91.11(1.48) 22.29(4.24) 25.96(5.05)
Hot Flip 83.26(5.87) 12.82(7.64) 84.18(2.25) 15.38(4.14) 23.07(3.67)
Random 62.57(6.63) 3.81(4.12) 74.06(0.75) 8.82(1.14) 12.29(3.33)
SCIDOCS NQ AGGD 21.05(7.78) 11.95(7.99) 9.78(1.97) 0.88(0.29) 0.38(0.08)
Hot Flip 27.02(13.4) 14.32(4.48) 7.1(0.87) 1.18(0.4) 0.35(0.11)
Random 12.88(3.76) 0.78(0.61) 3.05(0.93) 0.6(0.37) 0.3(0.07)
MS MARCO AGGD 23.22(6.99) 14.3(20.69) 31.9(7.2) 3.88(1.62) 1.08(0.42)
Hot Flip 24.98(6.33) 12.25(7.52) 21.88(4.33) 3.12(1.4) 0.85(0.32)
Random 15.0(1.53) 1.27(1.83) 10.62(5.11) 2.48(1.38) 0.2(0.07)
SciFact NQ AGGD 22.92(4.46) 1.0(0.62) 6.75(1.92) - -
Hot Flip 24.5(12.16) 0.92(0.72) 3.58(0.98) 0.17(0.29) -
Random 8.25(3.42) - 0.5(0.37) - -
MS MARCO AGGD 29.08(10.43) 2.67(4.43) 30.33(6.47) 2.42(0.6) 0.5(0.29)
Hot Flip 21.58(6.52) 1.25(1.11) 16.5(4.78) 2.0(0.24) 0.67(0.71)
Random 9.83(5.46) - 7.33(4.96) 1.17(0.96) 0.17(0.17)
@@
Table 2: Out-of-domain top-20 attack success rate with only 1 adversarial passage. Results are averaged over 4 random runs with standard deviation shown in parenthesis. The combinations of attack and source dataset that achieve the highest ASR on each target domain and retriever are highlighted in bold.
@@
Out-of-Domain Attack We found that the generated adversarial passages can transfer across different domains. In Table 2, we use adversarial passages generated from training set of NQ and MS MARCO and insert them into the corpora of retrieval tasks in other domains. We found that (1) Compared to adversarial passages generated from NQ dataset, those trained from MS MARCO generally perform better in out-of-domain attacks, possibly because MS MARCO contains more training data. (2) Contriever models are still the most vulnerable to corpus poisoning attacks. For example, inserting a single adversarial passage generated by AGGD into FiQA-2018 achieves a top-$k_{r}=20$ ASR of 86.54%, and inserting it into Quora can trick the model into returning the adversarial passage in the top-20 retrieved passage for 86.92% of the queries. (3) Quora are easier domains to attack. AGGD achieves over 20% ASR even in DPR-mul and DPR-nq, which is surprising since, as shown in the in-domain attack results in Table 1, attacking DPR-mul and DPR-nq is extremely hard , even with adversarial passage trained from in-domain data. (Though NFCorpus has high attack success rate, NFCorpus also has a smaller testing corpus).
@@ 5.3 Analysis and Ablation Study
The effect of the token length $m$
5.4 Extending AGGD to Knowledge Poisoning Attacks
@@ 6 Conclusion
7 Limitations and Future Work
While we use corpus poisoning attacks to showcase our AGGD algorithm, the proposed attack framework is versatile and applicable to a wide range of adversarial attack scenarios, as many of these can be formulated as discrete optimization problems. However, the generated output sequences often lack semantic coherence, making the adversarial corpus easily detectable and filterable. A promising direction for future work is to reformulate the problem as a constrained optimization problem, focusing on producing semantically meaningful adversarial passages that are more difficult to defend against, even if this may compromise the overall attack success rate.
8 Ethics Statement
Our work studies the vulnerability of dense retriever and corpus poison attack. The propose attack AGGD shows higher attach success rate especially for ANCE model, compared to previous HotFlip attack, which could be used for spread garbage information. Future research based on this paper should be exercised with caution and consider the potential consequences.